12. Under physiological conditions, the true free-energy change for the hydrolysis of ATP to ADP is influenced by a number of factors including the presence of an associated Mg2+ ion nd is about-50 kJ mol-1. Suppose an individual maintain:s a constant amount of 100 g MgATP2. 0 NII H2N 9 HO MgATP2- a. How much free energy is represented by this amount of MgATP2i? b. ATP is not as much for energy storage as for energy transduction; ATP is hydrolyzed and restored over and over. If an individual turns over 6000 kJ in energy via ATP hydrolysis, how many times per day is this 100 g of ATP replenished?
Answer #1

a) the equation is

MgATP2- +H2O ----> MgADP- + Pi

Molar mass of   MgATP2- is 528.5 g / mole

Given mass of MgATP2- is 100grams

so moles of MgATP2- will be

Moles = \frac{Mass}{Molar mass} = \frac{100}{528.5} = 0.189 moles

So amount of free energy represented by 100g = 0.189 X (-50kJ) = -9.45kJ

b) Number of times per day this 100g replenished = -6000 / -9.45kJ = 634.92

c) For 600kJ of energy number of moles required = 600 /50 = 12 moles

so mass required = 12 X 528.5 = 6342 grams